# Aeos: Transport of metals from minihalos following Population III stellar feedback

### v1

@@ Abstract
We investigate how stellar feedback from the first stars (Population III) distributes metals through the interstellar and intergalactic medium using the star-by-star cosmological hydrodynamics simulation, Aeos. We find that energy injected from the supernovae of the first stars is enough to expel a majority of gas and injected metals beyond the virial radius of halos with mass $M_{*}\lesssim 10^{7}$ M, regardless of the number of supernovae. This prevents self-enrichment and results in a non-monotonic increase in metallicity at early times. Most minihalos ($M\gtrsim 10^{5}\,\rm M_{\odot}$) do not retain significant fractions of the yields produced within their virial radii until they have grown to halo masses of $M\gtrsim 10^{7}\,\rm M_{\odot}$. The loss of metals to regions well beyond the virial radius delays the onset of enriched star formation and extends the period that Population III star formation can persist. We also explore the contributions of different nucleosynthetic channels to 10 individual elements. On the timescale of the simulation (lowest redshift $z=14.3$), enrichment is dominated by core-collapse supernovae for all elements, but with a significant contribution from asymptotic giant branch winds to the s-process elements, which are normally thought to only be important at late times. In this work, we establish important mechanisms for early chemical enrichment which allows us to apply Aeos in later epochs to trace the evolution of enrichment during the complete transition from Population III to Population II stars.
@@ 1 Introduction
At redshifts of $3020<z<$ 30, mere hundreds of millions of years after the Big Bang, the first stars (Population III, or Pop III) began lighting up the Universe, ending the Dark Ages. Their contribution to the chemodynamical evolution of the Universe as well as their imprint on observed stars today are two of the major frontiers in astrophysics (Frebel & Norris, 2015; Klessen & Glover, 2023).
@@
In $\Lambda$-Cold Dark Matter cosmology, Pop III stars are expected to form from metal-free primordial gas that collapses in dark matter minihalos with masses of $10^{5}$$10^{6}\rm{M}_{\odot}$ (e.g. Haiman et al., 1996b; Tegmark et al., 1997; Bromm et al., 1999; Abel et al., 2002; Yoshida et al., 2003; Kulkarni et al., 2021). The lack of metals in primordial gas makes H2 ro-vibrational lines the only viable process for cooling at low temperatures ($T<10^{4}$ K) (Haiman et al., 1996a; Tegmark et al., 1997), resulting in cooling times longer than the free-fall time and, thus, large Jeans masses. Pop III stars, then, are thought to have a top-heavy initial mass function. Many studies have suggested that Pop III stars were very massive, with masses $10^{2}$$10^{3}\,\rm M_{\odot}$ (Abel et al., 2002; Bromm et al., 2002; Bromm & Loeb, 2004; Yoshida et al., 2008; Hirano et al., 2015; Hosokawa et al., 2016). However, more recent work has demonstrated that fragmentation from turbulence (Latif et al., 2013), weak magnetic fields (Sharda et al., 2020), and gravitational instabilities in disks (Stacy et al., 2010; Clark et al., 2011; Greif et al., 2011, 2012), as well as suppression of accretion by UV radiation (Hosokawa et al., 2011, 2016; Hirano et al., 2014; Latif et al., 2022) can reduce the characteristic mass of Pop III stars to tens of solar masses or lower.
@@
Regardless of the precise shape of the initial mass function, Pop III stars were likely too massive to survive until present day. However, their contribution to galactic and chemical evolution did not disappear upon their deaths through supernova (SN) explosions; instead they seeded their surrounding gas with the first metals from which the next generation of stars formed. Among them were slightly metal-enriched low-mass stars (Population II, or Pop II) (Ritter et al., 2016; Chiaki et al., 2016, 2018; Chiaki & Wise, 2019; Magg et al., 2022; Chen et al., 2024) that survived to the present (Gnedin & Kravtsov, 2006; Tumlinson, 2010; Griffen et al., 2018). These stars have retained the chemical fingerprints of Pop III in their atmospheres, which we can observe. Searching for these rare metal-poor stars is a non-trivial pursuit, but many have been found in the stellar halo as well as across nearly all dwarf satellite galaxies including the ultra-faint dwarf galaxies (UFDs; Frebel & Norris, 2015). UFDs inefficiently formed stars until being quenched by reionization (Brown et al., 2014) and are the ideal cosmic labs in which to study these stars. Furthermore, UFDs serve as foundational building blocks of larger galaxies, and our own Milky Way’s stellar halo is likely comprised of their remnants. The chemical abundances of metal-poor stars in the UFDs and the stellar halo are promising ways to investigate early galaxy formation and chemical enrichment (e.g., Ji et al., 2015; Brauer et al., 2019), and stellar chemical abundances have been studied through high-resolution spectroscopy and wide-field spectroscopic surveys such as APOGEE (Majewski et al., 2017), GALAH (De Silva et al., 2015), LAMOST (Cui et al., 2012a, b), and Gaia RVS (Gaia Collaboration et al., 2016, 2023).
@@
A few stars have been found whose abundances suggest they have been enriched by a single progenitor SN (Ezzeddine et al., 2019; Skúladóttir et al., 2021; Yong et al., 2021; Xing et al., 2023; Skúladóttir et al., 2024; Ji et al., 2024). However, our understanding of the progenitors of these stars rest heavily on our assumptions about the explosion mechanisms and yields of Pop III SNe (e.g. Heger & Woosley, 2002, 2010) and the dilution and mixing of metals in the interstellar and intergalactic medium (ISM and IGM). Thus, having a theoretical understanding of the chemical evolution of the first galaxies, ignited by the first generations of stars, has become critical to moving forward with interpreting the history of our Galaxy.
@@
Cosmological hydrodynamics simulations aimed at studying the chemodynamical evolution of galaxies have been largely successful at reproducing galaxy-scale relations such as the mass-metallicity relation (Obreja et al., 2014; Ma et al., 2016; Davé et al., 2017; De Rossi et al., 2017; Hirai et al., 2017; Torrey et al., 2019; Agertz et al., 2020; Marszewski et al., 2024), metallicity distribution functions, and the evolution of individual chemical abundances (Marcolini et al., 2008; Revaz et al., 2009; Sawala et al., 2010; Revaz & Jablonka, 2012; Jeon et al., 2017; Hirai & Saitoh, 2017; Hirai et al., 2017). Cosmological simulations cannot directly model small-scale physical processes given the large dynamic range, instead relying on sub-grid models to include these phenomena (see Crain & van de Voort, 2023, and references therein). Among these is the mixing of chemically inhomogeneous particles and regions. Until recently, the assumption of homogeneous mixing was sufficient to reproduce observed global abundance trends across galaxies, but the influx of copious amounts of chemical abundance data for Milky Way and Local Group stars in the last decade has unveiled new variations in the chemical abundances of coeval and co-natal populations of stars (Hill et al., 2019; Ji et al., 2020; Mead et al., 2024) that can only be explained by more complex evolutionary processes.
@@
Recent computational advances have opened the door to more directly modeling smaller-scale physics in cosmological simulations by increasing both mass and spatial resolution, thereby reducing the dependence on sub-grid models. This has been a critical and necessary advance toward understanding the more complex galaxy formation processes that drive chemical evolution, particularly in the early universe and dwarf galaxies. These higher resolution simulations have begun to model single stars as individual particles (Emerick et al., 2019; Gutcke et al., 2021; Lahén et al., 2020, 2023; Andersson et al., 2023, 2024). This enables the study of individual stellar feedback and how the details of when and where individual stars inject energy and enriched material impact the chemodynamical evolution of their environment.
@@
In this paper, we use Aeos, a new star-by-star cosmological hydrodynamics simulation (Brauer et al., 2024, hereafter, B24), to study the chemical enrichment of minihalos following Pop III and the first Pop II SNe, and the transport of metals between halos and the IGM. Theory and simulations have demonstrated that dwarf galaxies can efficiently enrich their circumgalactic medium (CGM) and the IGM (Dekel & Silk, 1986; Mac Low & Ferrara, 1999; Fragile et al., 2004; Muratov et al., 2017), showing that outflows of enriched material are primarily dependent on the mass of the halo and the location of the SNe. Large halos and halos with spatially dispersed SNe are more likely to retain their metals. However, studies of these outflows of enriched material have thus far been limited to halos $M_{\rm vir}>10^{6}\,\rm M_{\odot}$ or low redshift ($z<4$). Furthermore, studies on the distribution of metals in and beyond host galaxies have been extended to individual elements from different nucleosynthetic origins in theoretical models (Krumholz & Ting, 2018) and in hydrodynamic simulations (Emerick et al., 2018). These works have found that the mixing and retention of individual elements in a galaxy are dependent on their astrophysical origin. With Aeos, we are now able to extend these studies to high redshift minihalos that are the progenitors to dwarf galaxies.
@@
This paper is organized as follows: in Section 2 we describe the Aeos model, but we refer the reader to B24 for additional details. We explore the behavior of metals in the multiphase ISM and IGM in Section 3 and characterize the loss of metals from halos to the IGM in Section 4. In Section 5, we take a look at the individual elements traced in Aeos, specifically the contributions from different nucleosynthetic sources, and their individual behaviors moving from the ISM to IGM. Finally, in Section 6, we summarize our findings and discuss the implications of our results for future work in numerical and analytical models.
@@ 2 Simulation
Aeos is a set of high-resolution cosmological simulations with a box size of 1 comoving Mpc per side using the same initial conditions as Skinner & Wise (2020) that treats star particles as individual stars sampled from an assumed IMF, or star-by-star, rather than as single stellar populations. The methods used in these simulations are described in B24, and for the case of an idealized galaxy, in Emerick et al. (2019, hereafter, E19). The description here is limited to a summary of the properties and physics relevant to this paper.
We follow non-equilibrium primordial chemistry using Grackle (Smith et al., 2017) with radiative heating and self-shielding from a Haardt & Madau (2012) UV background. To limit the number of radiation sources while still capturing the vast majority of the photon energy budget, we restrict radiation to massive stars with $M_{*}>$8 M. Ionizing radiation from massive stars is tracked using the Moray adaptive ray-tracing radiative transfer method (Wise & Abel, 2011). Additionally, we trace non-ionizing UV radiation from massive stars that causes photoelectric heating and Lyman-Werner (LW) dissociation of H2. We track stellar feedback from supernovae and stellar winds, and follow yields for 10 metal species (C, N, O, Na, Mg, Ca, Mn, Fe, Sr, Ba) in addition to the total metal density.
@@ 2.1 Hydrodynamics
Aeos uses the adaptive mesh refinement (AMR) hydrodynamics code Enzo (Bryan et al., 2014; Brummel-Smith et al., 2019) to simulate the chemodynamical evolution and feedback of the early universe in high-resolution, with a maximum grid refinement of 1 proper pc. At this resolution, the Jeans length can become unresolved, leading to artificial fragmentation. While in E19, a pressure floor was used to support gas against artificial fragmentation, Aeos does not use one, instead setting the star formation density sufficiently low to avoid under-resolving the Jeans length and relying on the efficient conversion of dense gas into stars.
@@ 2.2 Star Formation
Stars form stochastically if the following conditions are met by assuming that the local star formation rate is proportional to an efficiency per free-fall time, which we take as $\epsilon_{\rm ff}=0.02$. The conditions for star formation are: i) $\nabla\cdot v<0$ (converging gas flow), ii) $T<500$ K, and iii) $n>10^{4}$ cm-3.
@@ 2.2.1 Population III Star Formation
Pop III stellar masses are sampled from a Salpeter IMF (Salpeter, 1955) with a power-law slope $\alpha=2.3$ for masses above characteristic mass, $M_{\rm char}$, and an exponential cut-off below (following Wise et al., 2012). As mentioned in B24, there are two runs of the simulation which use different mass ranges and characteristic masses for Pop III. In this paper, we use the Aeos20 simulation which has a minimum Pop III mass of 1 $\rm M_{\odot}$ and maximum mass of 300 $\rm M_{\odot}$, with $M_{\rm char}$ = 20 $\rm M_{\odot}$. Thus, most of our Pop III stars will have masses between $20-300\rm M_{\odot}$. Pop III stars are assigned lifetimes from Schaerer (2002).

### v2

@@ Abstract
We investigate how stellar feedback from the first stars (Population III) distributes metals through the interstellar and intergalactic medium using the star-by-star cosmological hydrodynamics simulation, Aeos. We find that energy injected from the supernovae of the first stars is enough to expel a majority of gas and injected metals beyond the virial radius of halos with mass $M_{\rm dm}\lesssim 10^{7}$ M, regardless of the number of supernovae. This prevents self-enrichment and results in a nonmonotonic increase in metallicity at early times. Most minihalos ($M_{\rm dm}\gtrsim 10^{5}\,\rm M_{\odot}$) do not retain significant fractions of the yields produced within their virial radii until they have grown to halo masses of $M_{\rm dm}\gtrsim 10^{7}\,\rm M_{\odot}$. The loss of metals to regions well beyond the virial radius delays the onset of enriched star formation and extends the period that Population III star formation can persist. We also explore the contributions of different nucleosynthetic channels to 10 individual elements. On the timescale of the simulation (lowest redshift $z=14.3$), enrichment is dominated by core-collapse supernovae for all elements, but with a significant contribution from asymptotic giant branch winds to the s-process elements, which are normally thought to only be important at late times. In this work, we establish important mechanisms for early chemical enrichment, which allows us to apply Aeos in later epochs to trace the evolution of enrichment during the complete transition from Population III to Population II stars.
@@ 1 Introduction
At redshifts of $20<z<30$, mere hundreds of millions of years after the Big Bang, the first stars (Population III, or Pop III) began lighting up the Universe, ending the Dark Ages. Their contribution to the chemodynamical evolution of the Universe as well as their imprint on observed stars today are two of the major frontiers in astrophysics (Frebel & Norris, 2015; Klessen & Glover, 2023).
@@
In $\Lambda$-cold dark matter cosmology, Pop III stars are expected to form from metal-free primordial gas that collapses in dark matter minihalos with masses of $10^{5}$$10^{6}\rm{M}_{\odot}$ (e.g. Haiman et al., 1996b; Tegmark et al., 1997; Bromm et al., 1999; Abel et al., 2002; Yoshida et al., 2003; Kulkarni et al., 2021). The lack of metals in primordial gas makes H2 rovibrational lines the only viable process for cooling at low temperatures ($T<10^{4}$ K) (Haiman et al., 1996a; Tegmark et al., 1997), resulting in cooling times longer than the freefall time and, thus, large Jeans masses. Pop III stars, then, are thought to have a top-heavy initial mass function (IMF). Many studies have suggested that Pop III stars were very massive, with masses $10^{2}$$10^{3}\,\rm M_{\odot}$ (Abel et al., 2002; Bromm et al., 2002; Bromm & Loeb, 2004; Yoshida et al., 2008; Hirano et al., 2015; Hosokawa et al., 2016). However, more recent work has demonstrated that fragmentation from turbulence (Latif et al., 2013), weak magnetic fields (Sharda et al., 2020), and gravitational instabilities in disks (Stacy et al., 2010; Clark et al., 2011; Greif et al., 2011a, 2012), as well as suppression of accretion by UV radiation (Hosokawa et al., 2011, 2016; Hirano et al., 2014; Latif et al., 2022) can reduce the characteristic mass of Pop III stars to tens of solar masses or lower. A less-top-heavy IMF is also supported by current observational constraints from the numbers and properties of carbon-enhanced metal-poor stars at [Fe/H] <5<-5, which disfavor a flat Pop III IMF of $10-300\rm M_{\odot}$ (de Bennassuti et al., 2017).
Regardless of the precise shape of the IMF, Pop III stars were likely too massive to survive until present day. Although several theoretical studies have demonstrated the ability to form Pop III stars with masses below $0.8\,\rm M_{\odot}$ (e.g. Jaura et al., 2022; Sharda & Menon, 2025; Lake et al., 2025), the non-detection of Pop III stars in Local Group ultrafaint dwarf galaxies (UFDs) thus far seems to reject this possibility. Modeling from Hartwig et al. (2015) and Magg et al. (2019) has shown that given the non-detection of metal-free stars, the probability of Pop III survivors in our stellar halo is unlikely. Rossi et al. (2021) also demonstrate, by simulating the color-magnitude diagram of Boötes I and comparing to 96 stars with metallicity measurements, that the Pop III $m_{\rm min}>0.8\,\rm M_{\odot}$. Additionally, semi-analytic modeling from Hartwig et al. (2024) has found a best fit $M_{\rm min}=13.6\rm M_{\odot}$.
@@
Even if they do not survive until today, the contribution of Pop III stars to galactic and chemical evolution did not disappear upon their deaths through supernova (SN) explosions; instead, they seeded their surrounding gas, and in some cases, nearby halos (e.g. Smith et al., 2015), with the first metals from which the next generation of stars formed. Among them were slightly metal-enriched low-mass stars (Population II, or Pop II) (Chiaki et al., 2016, 2018; Chiaki & Wise, 2019; Magg et al., 2022; Chen et al., 2024) that survived to the present (Gnedin & Kravtsov, 2006; Tumlinson, 2010; Griffen et al., 2018). These stars have retained the chemical fingerprints of Pop III in their atmospheres, which we can observe. Searching for these rare metal-poor stars is a nontrivial pursuit, but many have been found in the stellar halo as well as across nearly all dwarf satellite galaxies including UFDs (Frebel & Norris, 2015). UFDs inefficiently formed stars until being quenched by reionization (Brown et al., 2014) and are the ideal cosmic labs in which to study these stars. Furthermore, UFDs serve as foundational building blocks of larger galaxies, and our own Milky Way’s stellar halo is likely comprised of their remnants. The chemical abundances of metal-poor stars in the UFDs and the stellar halo are promising ways to investigate early galaxy formation and chemical enrichment (e.g., Ji et al., 2015; Brauer et al., 2019), and stellar chemical abundances have been studied through high-resolution spectroscopy and wide-field spectroscopic surveys such as APOGEE (Majewski et al., 2017), GALAH (De Silva et al., 2015), LAMOST (Cui et al., 2012), and Gaia RVS (Gaia Collaboration et al., 2016, 2023).
@@
A few stars have been found whose abundances suggest they have been enriched by a single progenitor SN (Ezzeddine et al., 2019; Skúladóttir et al., 2021; Yong et al., 2021; Xing et al., 2023; Skúladóttir et al., 2024; Ji et al., 2024). However, our understanding of the progenitors of these stars rests heavily on our assumptions about the explosion mechanisms and yields of Pop III SNe (e.g. Heger & Woosley, 2002, 2010) and the dilution and mixing of metals in the interstellar medium (ISM) and intergalactic medium (IGM). Thus, having a theoretical understanding of the chemical evolution of the first galaxies, ignited by the first generations of stars, has become critical to moving forward with interpreting the history of our Galaxy.
@@
Cosmological hydrodynamics simulations aimed at studying the chemodynamical evolution of galaxies have been largely successful at reproducing galaxy-scale relations such as the mass-metallicity relation (Obreja et al., 2014; Ma et al., 2016; Davé et al., 2017; De Rossi et al., 2017; Hirai et al., 2017; Torrey et al., 2019; Agertz et al., 2020; Marszewski et al., 2024), metallicity distribution functions, and the evolution of individual chemical abundances (Marcolini et al., 2008; Revaz et al., 2009; Sawala et al., 2010; Revaz & Jablonka, 2012; Jeon et al., 2017; Hirai & Saitoh, 2017; Hirai et al., 2017). Cosmological simulations cannot directly model small-scale physical processes given the large dynamic range, instead relying on subgrid models to include these phenomena (see Crain & van de Voort, 2023, and references therein). Among these is the mixing of chemically inhomogeneous particles and regions. Until recently, the assumption of homogeneous mixing was sufficient to reproduce observed global abundance trends across galaxies, but the influx of copious amounts of chemical abundance data for Milky Way and Local Group stars in the last decade has unveiled new variations in the chemical abundances of coeval and conatal populations of stars (Hill et al., 2019; Ji et al., 2020; Mead et al., 2024) that can only be explained by more complex evolutionary processes.
@@
Recent computational advances have opened the door to more directly modeling smaller-scale physics in cosmological simulations by increasing both mass and spatial resolution, thereby reducing the dependence on sub-grid models. This has been a critical and necessary advance toward understanding the more complex galaxy formation processes that drive chemical evolution, particularly in the early Universe and dwarf galaxies. These higher-resolution simulations have begun to model single stars as individual particles (Emerick et al., 2019; Gutcke et al., 2021; Lahén et al., 2020, 2023; Andersson et al., 2023, 2025). This enables the study of individual stellar feedback and how the details of when and where individual stars inject energy and enriched material impact the chemodynamical evolution of their environment.
@@
In this paper, we use Aeos, a new star-by-star cosmological hydrodynamics simulation (Brauer et al., 2025, hereafter B24), to study the chemical enrichment of minihalos following Pop III and the first Pop II SNe, and the transport of metals between halos and the IGM. Theory and simulations have demonstrated that dwarf galaxies can efficiently enrich their circumgalactic medium and the IGM (Dekel & Silk, 1986; Mac Low & Ferrara, 1999; Fragile et al., 2004; Muratov et al., 2017), showing that outflows of enriched material are primarily dependent on the mass of the halo and the location of the SNe. Large halos and halos with spatially dispersed SNe are more likely to retain their metals. Other work has examined the effects of stellar feedback on star formation specifically in low-metallicity environments (e.g. Valentini et al., 2023). However, studies of these outflows of enriched material and the effect on subsequent star formation have thus far largely been limited to halos $M_{\rm vir}>10^{6}\,\rm M_{\odot}$ or low redshift ($z<4$). Furthermore, studies on the distribution of metals in and beyond host galaxies have been extended to individual elements from different nucleosynthetic origins in theoretical models (Krumholz & Ting, 2018) and in hydrodynamic simulations (Emerick et al., 2018). These works have found that the mixing and retention of individual elements in a galaxy are dependent on their astrophysical origin. With Aeos, we are now able to extend these studies to high redshift minihalos that are the progenitors to dwarf galaxies.
@@
This paper is organized as follows: In Section 2 we describe the Aeos model, but we refer the reader to B24 for additional details. We explore the behavior of metals in the multiphase ISM and IGM in Section 3 and characterize the loss of metals from halos to the IGM in Section 4. In Section 5, we take a look at the individual elements traced in Aeos, specifically the contributions from different nucleosynthetic sources, and their individual behaviors moving from the ISM to IGM. Finally, in Section 6, we summarize our findings, consider the effects of certain choices in our simulation on the robustness of our results, and discuss the implications of our results for future work in numerical and analytical models.
@@ 2 Simulation
Aeos is a set of high-resolution cosmological simulations with a box size of 1 comoving Mpc per side using the same initial conditions as Skinner & Wise (2020) that treat star particles as individual stars sampled from an assumed IMF, or star-by-star, rather than as single stellar populations. The methods used in these simulations are described in B24, and for the case of an idealized galaxy, in Emerick et al. (2019, hereafter, E19). The description here is limited to a summary of the properties and physics relevant to this paper.
We follow non-equilibrium primordial chemistry using Grackle (Smith et al., 2017) with radiative heating and self-shielding from a Haardt & Madau (2012) UV background. To limit the number of radiation sources while still capturing the vast majority of the photon energy budget, we restrict radiation to massive stars with $M_{*}>$8 M. Ionizing radiation from massive stars is tracked using the Moray adaptive ray-tracing radiative transfer method (Wise & Abel, 2011). Additionally, we trace non-ionizing UV radiation from massive stars that causes photoelectric heating and Lyman–Werner (LW) dissociation of H2. We track stellar feedback from SNe and stellar winds, and follow yields for 10 metal species (C, N, O, Na, Mg, Ca, Mn, Fe, Sr, Ba) in addition to the total metal density.
@@ 2.1 Hydrodynamics
Aeos uses the adaptive mesh refinement hydrodynamics code Enzo (Bryan et al., 2014; Brummel-Smith et al., 2019) to simulate the chemodynamical evolution and feedback of the early universe in high-resolution, with a maximum grid refinement of 1 proper pc. At this resolution, the Jeans length can become unresolved, leading to artificial fragmentation. While in E19, a pressure floor was used to support gas against artificial fragmentation; Aeos does not use one, instead setting the star formation density sufficiently low to avoid under-resolving the Jeans length and relying on the efficient conversion of dense gas into stars.
@@ 2.2 Star Formation
Stars form stochastically if the following conditions are met by assuming that the local star formation rate is proportional to an efficiency per freefall time, which we take as $\epsilon_{\rm ff}=0.02$. The conditions for star formation are: (i) $\nabla\cdot v<0$ (converging gas flow), (ii) $T<500$ K, and (iii) $n>10^{4}$ cm-3.
@@ 2.2.1 Population III Star Formation
Pop III stellar masses are sampled from a Salpeter IMF (Salpeter, 1955) with a power-law slope $\alpha=2.3$ for masses above characteristic mass, $M_{\rm char}$, and an exponential cutoff below (following Wise et al., 2012). As mentioned in B24, there are two runs of the simulation that use different mass ranges and characteristic masses for Pop III. In this paper, we use the Aeos20 simulation, which has a minimum Pop III mass of 1 $\rm M_{\odot}$ and maximum mass of 300 $\rm M_{\odot}$, with $M_{\rm char}$ = 20 $\rm M_{\odot}$. Thus, most of our Pop III stars will have masses between $20-300\rm M_{\odot}$. Pop III stars are assigned lifetimes from Schaerer (2002).